Can someone explain the math on how to solve, the answers are circled. Just want to know the way to solve to get those answers.Questions 2 - 3 refer to the following experiment: In a generalized transduction experiment donor E. coli cells have the genotype a+ c+ b-, and recipient cells have the genotype a- c b+. Pl-mediated transductants for a+ were selected, and their total genotypes were determined, with the following results: Genotype atctb Number of progen 205 atc-bt 195 500 2. What is the cotransduction percentage for a and b? A) 20% What is the cotransduction percentage for a and c? A) 200 B) 40% C) 58% D) 60 E) 80% 3. B) 40% C) 58% D) 60% E) 80%
Answer #1

Generally the co-transduction frequency is refers both quads and double crosses and these can be include with the donor form of the unselected marker for the given interval.

According to the given situation

a+cb- Progeny 95

a+cb- Progeny 205

a+cb+ Progeny 5

a+cb+ Progeny 195

                  Total 500

The gene order is a,b and c.

2. ANS: the cotransduction percent for a and b is 60%.

Explanation:

a b interval is 95+205/500= 300/500 = 0.6 or 60%.

(both 95 and 205 are associated with the b- genotype)

So the option D. 60% is correct.

3. ANS: the cotransduction percent for a and c is 20%.

Explanation:

a c interval is 95+ 5/500= 100/500 = 0.2 or 20%.

So the option A. 20% is correct.